# Closed-Form Evaluation of $\sum_{n=2}^{\infty}\operatorname{arctanh}(n^{-k})$ via Infinite Products

### v1

@@ 3 Proof of the main theorem
The second sum is divergent, so we proceed more carefully. We write
$\ln\!\left(\frac{f(k)}{g(k)}\right)=2\bigl[\zeta(k)-1\bigr]+2\sum_{n=1}^{\infty}\frac{\zeta\bigl(k(2n+1)\bigr)-1}{2n+1}+1+2\sum_{n=1}^{\infty}\frac{1}{2(2n)(2n+1)}.$ (16)
However, a cleaner route is the following. Starting from
$\ln\!\left(\frac{f(k)}{g(k)}\right)=2\bigl[\zeta(k)-1\bigr]+1+2\sum_{n=1}^{\infty}\frac{\zeta\bigl(k(2n+1)\bigr)-1}{2n+1}+2\sum_{n=1}^{\infty}\frac{1}{2n(2n+1)},$ (17)
we note that $\sum_{n=1}^{\infty}\frac{1}{2n(2n+1)}=\sum_{n=1}^{\infty}\bigl(\frac{1}{2n}-\frac{1}{2n+1}\bigr)=1-\ln 2$, a standard telescoping result. Hence
@@
$\ln\!\left(\frac{f(k)}{2\,g(k)}\right)=2\sum_{n=2}^{\infty}\operatorname{arctanh}\!\left(\frac{1}{n^{k}}\right),$
@@ 8.5 OEIS sequence data
9 Acknowledgment
The author thanks the anonymous referees for helpful comments.

### v2

@@ 3 Proof of the main theorem
Because the second sum is divergent, we regroup in adjacent odd-even blocks without reordering terms to keep all sums convergent:
$\ln\!\left(\frac{f(k)}{g(k)}\right)=2\zeta(k)-1-\sum_{n=1}^{\infty}\left[\frac{1}{2n}+\frac{1-2\zeta\!\bigl(k(2n+1)\bigr)}{2n+1}\right].$
Now, for the sum portion,
$\sum_{n=1}^{\infty}\left[\frac{1}{2n}+\frac{1-2\zeta\!\bigl(k(2n+1)\bigr)}{2n+1}\right]$ $=\sum_{n=1}^{\infty}\left[\frac{2n+1}{(2n)(2n+1)}+\frac{2n\!\left(1-2\zeta\!\bigl(k(2n+1)\bigr)\right)}{(2n)(2n+1)}\right]$
$=\sum_{n=1}^{\infty}\frac{1}{2n(2n+1)}+2\sum_{n=1}^{\infty}\frac{1-\zeta\!\bigl(k(2n+1)\bigr)}{2n+1}.$
Also,
$\sum_{n=1}^{\infty}\frac{1}{2n(2n+1)}=\sum_{n=1}^{\infty}\left(\frac{1}{2n}-\frac{1}{2n+1}\right)=1-\ln 2.$
Hence
@@
$\frac{1}{2}\ln\!\left(\frac{f(k)}{2\,g(k)}\right)=\sum_{n=2}^{\infty}\operatorname{arctanh}\!\left(\frac{1}{n^{k}}\right),$